Blood glucose levels for obese patients have a mean of 100 with a standard deviation of 15. A researcher thinks that a diet high in raw cornstarch will have a positive or negative effect on blood glucose levels. A Sample of 30 patients who have tried the raw cornstarch diet has a mean glucose level of 140.
Test the hypothesis that the raw cornstarch had an effect.
1
Expert's answer
2020-09-15T17:49:55-0400

Population standard deviation is known and the sample is large, "\ge30" , thus, normal distribution is used.

"H_o\mu=100"

"H_a\mu\ne100"

Assume 95% confidence level.

"CV, Z_{0.025}=1.96"

Test statistic"=\frac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}"

"=\frac{140-100}{\frac{15}{\sqrt{30}}}=14.61"

Since the calculated statistic 14.61 is grater than the critical value, 1.96, we reject the null hypothesis and conclude that the raw cornstarch had an effect on blood glucose level.